Bruce has $5.15 in dimes and quarters. He has 8 more quarters than dimes. How many quarters does he have.
Answer by mananth(16949)

let dimes be x
quarters = x+8
..
10x+25(x+8)=515
10x+25x+200=515
35x+200=515
-80
35x=315
/35
x=9 the number of dimes
quarters = 9+8 =17
...
check
10*9 +17*25=90+425=515
...
m.ananth@hotmail.ca